# Dynamic optical coherence tomography algorithm for label-free assessment of swiftness and occupancy of intratissue moving scatterers Supplementary Material

### v2

@@
Dynamic optical coherence tomography algorithm for label-free assessment of swiftness and occupancy of intratissue moving scatterers
Supplementary Material
††journal: opticajournal††articletype: Research ArticleDynamic optical coherence tomography (DOCT) statistically analyzes fluctuations in time-sequential OCT signals, enabling label-free and three-dimensional visualization of intratissue and intracellular activities. Current DOCT methods, such as logarithmic intensity variance (LIV) and OCT correlation decay speed (OCDS) have several limitations. Namely, the DOCT values and intratissue motions are not directly related, and hence DOCT values are not interpretable in the context of the tissue motility. We introduce a new open-source DOCT algorithm that provides more direct interpretation of DOCT in the contexts of dynamic scatterer ratio and scatterer speed in the tissue. The detailed properties of the new and conventional DOCT methods are investigated by numerical simulations based on our open-source DOCT simulation framework, and the experimental validation with in vitro and ex vivo samples demonstrates the feasibility of the new method.
@@ 1 Introduction
Several signal analysis algorithms have been developed for DOCT. One type of algorithm computes the variance or standard deviation of the time-sequential OCT signals, directly [14, 15, 12, 16, 17, 18] or indirectly [19, 20] This approach uses the magnitude of the OCT signal fluctuations to contrast intratissue and intracellular activities. Another type of algorithm computes the temporal autocorrelation of time-sequential OCT signals [15, 12, 17], and uses the decorrelation properties as the contrast source. Yet another type of algorithms analyzes the power spectral density of time-sequential OCT signals through a numerical Fourier transform [21, 22, 23, 24, 25, 26], and using the dominant frequency or powers within several frequency bins as the contrast source.
The present authors have previously developed two DOCT algorithms. The first is the logarithmic intensity variance (LIV), which is defined as the time variance of dB-scaled OCT signal intensity [17]. LIV quantifies the magnitude of signal scintillation, which had been considered to related the magnitude of intratissue activities. The other is the OCT correlation decay speed (OCDS), which is defined as the slope of the auto-correlation curve of the time-sequential OCT signal within a particular range of the auto-correlation delay time [17]. OCDS reflects the speed of OCT signal scintillation, which had been considered to relate to the speed of the intratissue activities. Previous studies have examined delay-time ranges of [12.8 ms, 64 ms] and [204.8 ms, 1228.8 ms], with the corresponding OCDS referred to as early OCDS (OCDSe) and late OCDS (OCDSl), respectively [17, 27]. OCDSe and OCDSl are considered to be indicators of fast and slow intratissue activities, respectively.
@@
Despite their feasibility and utility, the LIV and OCDS algorithms have two shortcomings. First, OCDS is not a direct indicator of the speed of dynamics. The selection of the delay-time range is arbitrary, and the correlation decay slope within a specific delay range neither directly nor monotonically correlates with the speed of the dynamics. Second, these DOCT methods quantify the dynamics of OCT signals rather than the dynamics of tissues. The relationship between the OCT- signal dynamics and tissue dynamics remains unclear. These two shortcomings hinder the interpretation of LIV and OCDS images in the context of tissue activities.
@@ 3.1.1 Dynamic sample modeling
The intratissue scatterers were assumed to follow not Brownian motion (i.e., diffusion) but random ballistic motion in which all scatterers move rectilinearly with the same constant speed , but with a motion direction that is randomly chosen. This selection of random ballistic motion was based on an extensive literature survey about the intratissue and intracellular activities[39]. Further points about the motion model will be discussed in Section 5.5. Assuming the initial position of the -th scatterer is (, ), the three-dimensional (3D) position of the scatterer after a particular travel time is given by
@@ 3.1.2 DOCT signal modeling
where is the scatterer index and is the number of scatterers in the sample. The term and the first exponential part (green part) correspond to the complex reflectivity of the -th scatterer. Here, is a coefficient accounting for the different reflectivity values of the scatterers, and we assume that this is 1 for all scatterers. The exponential part represents the phase defined by the depth position of the scatterer, where is the center wavelength of the probe beam and is the path length offset, defined by the arm lengths of the interference. is a constant for all scatterer contributions, and therefore it does not affect the intensity of the OCT signal.
@@ 3.3.2 Study 2: DSR dependency
Both aLIV and LIV increase monotonically with increasing DSR [Fig. 3(b, d)] when the scatterer speed is greater than 0.2 m/s. However, when the speed of the dynamic scatterers is less than 0.1 m/s, aLIV diverges to extremely large values. For speeds slower than 0.05 m/s, almost all aLIV values were plotted outside of the plot region. In contrast, LIV become smaller with decreasing scatterer speed, as shown in Fig. 3(d).
@@ 3.4 Summary of DOCT metrics’ properties
Swiftness increases almost monotonically as the scatterer speed increases, especially for speeds of less than 1.2 m/s. In addition, Swiftness is relatively insensitive to DSR. Thus, Swiftness can be used as an indicator for the speed of dynamic scatterers. It is noteworthy that, although the monotonic relation can be found only up to 1.2 m/s, this maximum speed of the monotonic relation can be enlarged by using a shorter interval of OCT frame acquisition.
OCDS displays a distinct peak at a particular scatterer speed, indicating that this metric can be used as a fingerprint for a specific scatterer speed. The principle of OCDS suggests that the fingerprint speed zone can be selected by changing the delay-time range. Furthermore, the location and width of the fingerprint speed zone can be identified from numerical simulations.
@@ 4.1.3 Study design and protocol
In Study 2, we investigated the impact of on the DOCT metrics and the image appearances. To examine the effect of , we virtually shortened by truncating the frame sequence, which initially contained 32 frames. Here, the virtually shortened is referred to as the virtual acquisition time window (). Six different were investigated, i.e., = 0.41, 1.23, 2.46, 3.69, 4.91, and 6.35 s. The time-sequential OCT data for each consisted of a different number of frames, i.e., 3, 7, 13, 19, 25, and 32 frames.
@@
It should be noted that the protocol of Study 2 is suboptimal, as the number of frames varies among the configurations. However, the effect of varying frame numbers is negligible for the following reasons. Specifically, all four types of DOCT metrics are derived from the variance of OCT signals. OCDS, in particular, is calculated from the correlation coefficient, which is defined using variances and covariance. As is well known in statistical theory, the expected value of variance is not significantly affected by the number of data points.
@@ 4.2.1 Study 1
Figure 4 presents the en face aLIV, LIV, Swiftness, OCDS, and fluorescence images (from left to right) of the untreated (1st row) and anti-cancer (1-M PTX) drug-treated tumor spheroids (treatment for 1, 3, and 6 days shown in 2nd, 3rd, and 4th rows, respectively). The fluorescence images were captured using a wide-field fluorescence microscope (THUNDER imager DMi8; Leica Micro-systems, Wetzer, Germany) with a microscopic objective lens (numerical aperture of 0.12). Living cells were highlighted with calcein-acetoxymethyl (calcein-AM; Dojindo, Kumamoto, Japan), which emits a green fluorescence signal, while dead cells were highlighted with propidium iodide (PI; Dojindo), which emits a red fluorescence signal.
For the without-drug and 1-day-treated spheroids, a concentric pattern consisting of two zones is observed in both the DOCT and fluorescence images [Fig. 4(a–j)]. The fluorescence images show dead cells (red) concentrated in the inner zone, while living cells (green) are distributed in the periphery. These characteristics are consistent with the well-known necrotic core formation of MCF-7 spheroids [47], which is caused by hypoxia and nutrient deficiency in the central region. Therefore, the inner and outer zones of the concentric pattern in the DOCT images may also correspond to the necrotic core and the living cells, respectively.
@@
For the 3-day-treated spheroid, the Swiftness and OCDS images reveal a clear three-zone concentric pattern [Fig. 4(m, n)]. Similarly, aLIV and LIV exhibit a concentric pattern of low (red)–moderate (yellow with red dots)–high (green) zones from the inner to outer regions [Fig. 4(k, l)]. In the innermost zone, a very small proportion of scatterers are dynamic (low aLIV and LIV) and move at high speed (high Swiftness). In the middle zone, both static and dynamic scatterers coexist (moderate aLIV and LIV), with dynamic scatterers moving at slower speeds (low Swiftness) within the fingerprint speed zone of OCDS (high OCDS). The outermost zone of the aLIV and LIV images resembles the outer zone of the without-drug and 1-day-treated spheroids, namely, a large proportion of the scatterers are dynamic (high aLIV and LIV). However, Swiftness and OCDS exhibit reverse contrasts to those of the without-drug and 1-day-treated spheroids. This indicates that the dynamic scatterers move rapidly (high Swiftness) and their speed is outside the fingerprint speed zone of OCDS (low OCDS). It is noteworthy that the fluorescence image does not differentiate the three concentric zones because the living and dead cells are intermixed.
@@
Vertical stripes appear in the B-scan images [Fig. 5(e–h)]. Mukherjee et al.investigated the LIV of mouse kidneys, and hypothesized that the vertical stripes of high LIV are artifacts caused by fast motion at the upper-most part of the stripes [33]. However, this hypothesis was not confirmed in that study because only LIV was used for DOCT imaging in Ref. [33]. On the other hand, the Swiftness image obtained in the present study provides supporting evidence. Namely, at the top of the vertical stripes [dotted circles in Fig. 5(i, j), which are magnified images of the square regions in Fig. 5(f, g)], high Swiftness values are observed, indicating high-speed dynamic scatterers. These regions exhibit low OCDS, probably because the scatterer speeds exceed the fingerprint speed zone of OCDS.
@@ 4.2.2 Study 2
The contrast of each DOCT metric between the core and periphery ROIs is plotted in Fig. 6(d). The OCDS of the core ROI exhibits erroneous negative values for small values, thus, the contrast of the OCDS was not computed. These erroneous values may be caused by the limited number of time points involved in the OCDS computation. For aLIV, LIV, and Swiftness, the contrast becomes large and stable as increases. It appears that aLIV stabilize faster than LIV. Considering that the DOCT values are more reliable at larger , aLIV might be a more reliable metric than LIV.
The contrast between the core and periphery ROIs for aLIV, LIV, and Swiftness decreases monotonically as becomes smaller [Fig. 6 (d)]. The aLIV metric demonstrates higher contrast than LIV, except at = 0.41 s. This higher contrast of aLIV agrees with our expectations based on the principle of aLIV. Namely, the saturation speed of the LIV curve is slow and it does no reach the saturation level within , such as in the case of the periphery ROI, the aLIV value, which is the expected saturation level of the curve, becomes larger than the LIV values. On the other hand, if the LIV curve saturates fast, such as in the case of core ROI, the aLIV and LIV values become close to each other. And hence, the contrast of aLIV between two ROIs can be higher than that of LIV, especially is small. Note that the case of the shortest (0.41 s) is an exception, where the is too short to correctly estimate the saturation level (that is the aLIV).
Figure 7 shows B-scan images of aLIV, LIV, Swiftness, and OCDS of a tumor spheroid at each . The observational contrasts are consistent with the contrast plot of Fig. 6. The image quality of all DOCT metrics degrades as decreases. According to subjective observations, the image quality of aLIV and LIV with = 3.69 s or larger is acceptable. In the Swiftness image, the spheroid core is distinguishable for = 1.23 s or larger, while for OCDS, the core is distinguishable for = 2.46 s or larger.
@@ 5.1 DOCT signals and intratissue/intracellular motions
Most of the DOCT algorithms are not direct measures of sample dynamics (i.e., intratissue and intracellular motions) but measures of OCT signal fluctuations. In general, complex OCT signal is expressed as a summation of phasor contributions from scatterers in the tissue. The time frequency of the complex OCT sequence may directly correspond to the vibration of the phasors, but the vibration frequency of a phasor does not necessarily equal the vibration frequency of the scatterers. It can be easily imagined that the phase of the phasor rotates even when the scatterer moves with constant linear motion. Furthermore, several Fourier-spectrum-based DOCT methods use the time-frequency spectrum of the intensity or amplitude OCT signal sequence [21, 22, 23, 24, 25, 26]. According to discrete-scatterer-based formulations of OCT [48, 35], the intensity of OCT consists of interaction terms (i.e., the multiplications) of the phasors (see Section 3.2 and Fig. 2 of Ref. [35]). These interactions generate new frequency components that were not in the original complex OCT sequence. And hence, the Fourier-spectrum-based DOCT contrasts are not really direct measures of the sample motion. Note that, since OCT amplitude is defined as the square root of intensity, amplitude-based methods are also not free from this issue. (Note that the Taylor series expansion of the square root may have infinitely high order terms.)
Similarly, autocorrelation-based methods [15, 12, 17] are also not really direct measures of sample dynamics. It can be understood from the fact that the autocorrelation function and the power spectrum are tightly related by Wiener-Khinchin theorem. Note that, among the autocorrelation-based methods, OCDS [17] uses logarithmic intensity, and the Taylor series expansion of the logarithmic function also consists infinitely high order terms. And hence, OCDS is also not a direct measure of the sample dynamics.
Our new DOCT metrics, i.e., aLIV and Swiftness, are also not free from this issue. However, the numerical analyses revealed relatively straightforward relationships between these metrics and the sample dynamics (Section 3). Namely, aLIV mainly corresponds to DSR, while Swiftness mainly corresponds to the scatterer speed. These simple relationships enable easy interpretation of DOCT images. In addition, these properties of aLIV and Swiftness will also enable quantitative measurement of DSR and scatterer speed by combining with a sophisticated estimation theory as shown in a preliminary demonstration[49].

### v3

# Supplementary Material: Supplement 1

@@
Supplementary Material: Supplement 1
††journal: opticajournal††articletype: Research ArticleDynamic optical coherence tomography (DOCT) statistically analyzes fluctuations in time-sequential OCT signals, enabling label-free and three-dimensional visualization of intratissue and intracellular activities. Current DOCT methods, such as logarithmic intensity variance (LIV) and OCT correlation decay speed (OCDS) have several limitations. Namely, the DOCT values and intratissue motions are not directly related, and hence DOCT values are not interpretable in the context of the tissue motility. We introduce an open-source DOCT algorithm that provides more direct interpretation of DOCT in the contexts of dynamic scatterer ratio and scatterer speed in the tissue. The detailed properties of the new and conventional DOCT methods are investigated by numerical simulations based on our open-source DOCT simulation framework, and the experimental validation with in vitro and ex vivo samples demonstrates the feasibility of the method.
@@ 1 Introduction
Several signal analysis algorithms have been developed for DOCT. One type of algorithm computes the variance or standard deviation of the time-sequential OCT signals, directly [14, 15, 12, 16, 17, 18] or indirectly [19, 20]. This approach uses the magnitude of the OCT signal fluctuations to contrast intratissue and intracellular activities. Another type of algorithm computes the temporal autocorrelation of time-sequential OCT signals [15, 12, 17], and uses the decorrelation properties as the contrast source. Yet another type of algorithm analyzes the power spectral density of time-sequential OCT signals through a numerical Fourier transform [21, 22, 23, 24, 25, 26, 27], and uses the dominant frequency or powers within several frequency bins as the contrast source. The final type of algorithm analyzes the phase difference between the complex OCT signals at two adjacent time points [28], and uses the average and standard deviation of all the phase differences as the contrast source.
The present authors have previously developed two DOCT algorithms. The first is the logarithmic intensity variance (LIV), which is defined as the time variance of dB-scaled OCT signal intensity [17]. LIV quantifies the magnitude of signal scintillation, which had been considered to be related to the magnitude of intratissue activities. The other is the OCT correlation decay speed (OCDS), which is defined as the slope of the auto-correlation curve of the time-sequential OCT signal within a particular range of the auto-correlation delay time [17]. OCDS reflects the speed of OCT signal scintillation, which had been considered to be related to the speed of the intratissue activities. Previous studies have examined delay-time ranges of [12.8 ms, 64 ms] and [204.8 ms, 1228.8 ms], with the corresponding OCDS referred to as early OCDS (OCDSe) and late OCDS (OCDSl), respectively [17, 29]. OCDSe and OCDSl are considered to be indicators of fast and slow intratissue activities, respectively.
@@
Despite their feasibility and utility, the LIV and OCDS algorithms have two shortcomings. First, OCDS is not a direct indicator of the speed of dynamics. The selection of the delay-time range is arbitrary, and the correlation decay slope within a specific delay range neither directly nor monotonically correlates with the speed of the dynamics. For example, in drug testing, the responses of in vitro samples to various drug types and doses are evaluated. This requires a modality capable of sensitively distinguishing between different types of cellular activities. We believe that biological conditions can be effectively differentiated by identifying the speed of intratissue dynamics. Similar to OCDS, frequency-based DOCT methods generate contrast values relating to the fluctuation speed of OCT signals. However, the frequency of OCT signal is not a direct measure of the motion speed of the intratissue dynamic. Second, these DOCT methods quantify the dynamics of OCT signals rather than the dynamics of tissues. The relationship between the OCT-signal dynamics and tissue dynamics remains unclear. These two shortcomings hinder the interpretation of LIV and OCDS images in the context of tissue activities.
@@ 3.1.1 Dynamic sample modeling
The intratissue scatterers were assumed to follow not Brownian motion (i.e., diffusion) but random ballistic motion in which all scatterers move rectilinearly with the same constant speed , but with a motion direction that is randomly chosen. This selection of random ballistic motion was based on an extensive literature survey about the intratissue and intracellular activities and simulation results using the motion models [41]. Further points about the motion model will be discussed in Section 5.5. Assuming the initial position of the -th scatterer is (, ), the three-dimensional (3D) position of the scatterer after a particular travel time is given by
@@ 3.1.2 DOCT signal modeling
where is the scatterer index and is the number of scatterers in the sample. The term and the first exponential part (blue part) correspond to the complex reflectivity of the -th scatterer. Here, is a coefficient accounting for the different reflectivity values of the scatterers, and we assume that this is 1 for all scatterers. The exponential part represents the phase defined by the depth position of the scatterer, where is the center wavelength of the probe beam and is the path length offset, defined by the arm lengths of the interference. is a constant for all scatterer contributions, and therefore it does not affect the intensity of the OCT signal.
@@ 3.3.2 Study 2: DSR dependency
Both aLIV and LIV increase monotonically with increasing DSR [Fig. 3(b, d)] when the scatterer speed is greater than 0.2 m/s. However, when the speed of the dynamic scatterers is less than 0.1 m/s, aLIV diverges to extremely large values. For speeds slower than 0.05 m/s, almost all aLIV values were plotted outside of the plot region. In contrast, LIV becomes smaller with decreasing scatterer speed, as shown in Fig. 3(d).
@@ 3.4 Summary of DOCT metrics’ properties
Swiftness increases almost monotonically as the scatterer speed increases, especially for speeds of less than 1.2 m/s. In addition, Swiftness is relatively insensitive to DSR. Thus, Swiftness can be used as an indicator for the speed of dynamic scatterers. It is noteworthy that, although the monotonic relation can be found only up to 1.2 m/s, this maximum speed of the monotonic relation can be controlled by the OCT system speed and the scanning protocol. Specifically, higher speeds can be measured by using a shorter interval of OCT frame acquisition.
OCDS displays a distinct peak at a particular scatterer speed, indicating that this metric can be used as a fingerprint for a specific scatterer speed. Namely, the high-value of OCDS indicates the scatterer motion at a specific speed range. The principle of OCDS suggests that the fingerprint speed zone can be selected by changing the delay-time range. Namely, if we double the inter-frame time for example, the finger-print speed also doubles. Furthermore, the location and width of the fingerprint speed zone can be identified from numerical simulations. It should be noted that, in contrast to the markedness of the high OCDS value, the low OCDS cannot discriminate whether the scatterer speed is slower or faster than the specific range.
@@ 4.1.3 Study design and protocol
In Study 2, we investigated the impact of on the DOCT metrics and the image appearances. It is because low dependency of DOCT metrics on can be beneficial for future applications that require shorter acquisition times, such as drug screening by large numbers of in vitro samples and in vivo imaging. Shorter acquisition times enable longitudinal assessment of in vitro samples at fine measurement time intervals and help reduce the effects of bulk motion during in vivo imaging.
To examine the effect of , we virtually shortened by truncating the frame sequence, which initially contained 32 frames. Here, the virtually shortened is referred to as the virtual acquisition time window (). Six different were investigated, i.e., = 0.41, 1.23, 2.46, 3.69, 4.91, and 6.35 s. The time-sequential OCT data for each consisted of a different number of frames, i.e., 3, 7, 13, 19, 25, and 32 frames.
@@
It should be noted that the protocol of Study 2 is suboptimal, as the number of frames varies among the configurations. However, the effect of varying frame numbers is expected to be small for the following reasons. Specifically, all four types of DOCT metrics are derived from the variance of OCT signals. OCDS, in particular, is calculated from the correlation coefficient, which is defined using variances and covariance. As is well known in statistical theory, the stability of the variance estimate decreases as the number of data points decreases. This reduction in stability will reduce the overall image quality of the DOCT images. On the other hand, the expectation of variance is not significantly affected by the number of data points.
@@ 4.2.1 Study 1
Figure 4 presents the en face aLIV, LIV, Swiftness, OCDS, and fluorescence images (from left to right) of the untreated (1st row) and anti-cancer (1-M PTX) drug-treated tumor spheroids (treatment for 1, 3, and 6 days shown in 2nd, 3rd, and 4th rows, respectively). The fluorescence images were captured using a wide-field non-confocal fluorescence microscope (THUNDER imager DMi8; Leica Micro-systems, Wetzer, Germany) with a microscopic objective lens having a numerical aperture (NA) of 0.12. Living cells were highlighted with calcein-acetoxymethyl (calcein-AM; Dojindo, Kumamoto, Japan), which emits a green fluorescence signal, while dead cells were highlighted with propidium iodide (PI; Dojindo), which emits a red fluorescence signal. Since the NA of the fluorescence imaging is low, the fluorescence image is not really depth sectioned. This should be taken into account when the en face slice of DOCT and the fluorescence image are compared.
For the without-drug and 1-day-treated spheroids, a concentric pattern consisting of two zones is observed in both the DOCT and fluorescence images [Fig. 4(a-j)]. The fluorescence images show dead cells (red) concentrated in the inner zone, while living cells (green) are distributed in the periphery. These characteristics are consistent with the well-known necrotic core formation of MCF-7 spheroids [50], which is caused by hypoxia and nutrient deficiency in the central region. Therefore, the inner and outer zones of the concentric pattern in the DOCT images may also correspond to the necrotic core and the living cells, respectively.
Four representative LIV curves (curves 1-4) and their fitting curves are shown in Fig. 4(u). The curves 1 and 2 are obtained at the inner zone, while curves 3 and 4 are obtained at the outer zone as indicated in Fig. 4(v, w). These two zones exhibited distinct characteristics in their LIV curves. Specifically, the outer zone [blue plots in Fig. 4(u)] shows the LIV curves with a larger magnitude than the inner zone (red plots), while the inner zone shows the LIV curve with a faster saturation than the outer zone
@@
For the 3-day-treated spheroid, the Swiftness and OCDS images reveal a clear three-zone concentric pattern [Fig. 4(m, n)]. Similarly, aLIV and LIV exhibit a concentric pattern of low (red)–moderate (yellow with red dots)–high (green) zones from the inner to outer regions [Fig. 4(k, l)]. In the innermost zone, a very small proportion of scatterers are dynamic (low aLIV and LIV) and move at high speed (high Swiftness). In the middle zone, both static and dynamic scatterers coexist (moderate aLIV and LIV), with dynamic scatterers moving at slower speeds (low Swiftness) within the fingerprint speed zone of OCDS (high OCDS). The outermost zone of the aLIV and LIV images resembles the outer zone of the without-drug and 1-day-treated spheroids, namely, a large proportion of the scatterers are dynamic (high aLIV and LIV). However, Swiftness and OCDS exhibit reverse contrasts to those of the without-drug and 1-day-treated spheroids. This indicates that the dynamic scatterers move rapidly (high Swiftness) and their speed is outside the fingerprint speed zone of OCDS (low OCDS). It is noteworthy that the three concentric zones were not visible in the fluorescence image, which visualizes a mixture of live and dead cells.
@@
Vertical stripes appear in the B-scan images [Fig. 5(e–h)]. Mukherjee et al.investigated the LIV of mouse kidneys, and hypothesized that the vertical stripes of high LIV are artifacts caused by fast motion at the upper-most part of the stripes, and hence this vertical stripe is similar to projection artifacts observed in OCT angiography of the retina [35, 42]. However, this hypothesis was not confirmed in that study because only LIV was used for DOCT imaging in Ref. [35]. On the other hand, the Swiftness image obtained in the present study provides supporting evidence. Namely, at the top of the vertical stripes [dotted circles in Fig. 5(i, j), which are magnified images of the square regions in Fig. 5(f, g)], high Swiftness values are observed, indicating high-speed dynamic scatterers. These regions exhibit low OCDS, probably because the scatterer speeds exceed the fingerprint speed zone of OCDS.
It should be noted that the interpretations according to the simulation results rely on several assumptions, such as that a single motion model (i.e., random ballistic motion model) and a single scatterer speed govern the motion. The scatterer speed used in the simulation is confined to a reasonable range for the scatterers in the samples. These limitations are discussed in detail in the discussion section (Section 5.5).
@@ 4.2.2 Study 2
The contrast of each DOCT metric between the core and periphery ROIs is plotted in Fig. 6(d). The OCDS of the core ROI exhibits erroneous negative values for small values, thus, the contrast of the OCDS was not computed. These erroneous values may be caused by the limited number of time points involved in the OCDS computation. For aLIV, LIV, and Swiftness, the contrast becomes large and stable as increases. It appears that aLIV stabilizes faster than LIV. Considering that the DOCT values are more reliable at larger , aLIV might be a more reliable metric than LIV.
The contrast between the core and periphery ROIs for aLIV, LIV, and Swiftness decreases monotonically as becomes smaller [Fig. 6 (d)]. The aLIV metric demonstrates higher contrast than LIV, except at = 0.41 s. This higher contrast of aLIV agrees with our expectations based on the principle of aLIV. Namely, the saturation speed of the LIV curve is slow and it does not reach the saturation level within , such as in the case of the periphery ROI, the aLIV value, which is the expected saturation level of the curve, becomes larger than the LIV values. On the other hand, if the LIV curve saturates fast, such as in the case of core ROI, the aLIV and LIV values become close to each other. And hence, the contrast of aLIV between two ROIs can be higher than that of LIV, especially is small. Note that the case of the shortest (0.41 s) is an exception, where the is too short to correctly estimate the saturation level (that is the aLIV).
Figure 7 shows B-scan images of aLIV, LIV, Swiftness, and OCDS of a tumor spheroid at each . The observational contrasts are consistent with the contrast plot of Fig. 6. The image quality of all DOCT metrics degrades as decreases. According to subjective observations, the image quality of aLIV and LIV with = 3.69 s or larger is acceptable. In the Swiftness image, the spheroid core is distinguishable for = 1.23 s or larger, while for OCDS, the core is distinguishable for = 2.46 s or larger. The degradation of the image quality at regions with small values might be caused by the instability of the variance estimate due to the reduced number of data points.
@@ 5.1 DOCT signals and intratissue/intracellular motions
Most of the DOCT algorithms are not direct measures of sample dynamics (i.e., intratissue and intracellular motions) but measures of OCT signal fluctuations. In general, complex OCT signal is expressed as a summation of phasor contributions from scatterers in the tissue. The time frequency of the complex OCT sequence, e.g., obtained by a Fourier-spectrum-based DOCT method [51], may directly correspond to the vibration of the phasors, but the vibration of a phasor does not necessarily indicate the vibration of the scatterers. It can be easily imagined that the phase of the phasor rotates even when the scatterer moves with constant linear motion. Furthermore, several Fourier-spectrum-based DOCT methods use the time-frequency spectrum of the intensity or amplitude OCT signal sequence [21, 22, 23, 24, 25, 26]. According to discrete-scatterer-based formulations of OCT [52, 37], the intensity of OCT consists of interaction terms (i.e., the multiplications) of the phasors (see Section 3.2 and Fig. 2 of Ref. [37]). These interactions generate new frequency components that were not in the original complex OCT sequence. And hence, the Fourier-spectrum-based DOCT contrasts are not really direct measures of the sample motion. Note that, since OCT amplitude is defined as the square root of intensity, amplitude-based methods are also not free from this issue. (Note that the Taylor series expansion of the square root may have infinitely high order terms.)
Similarly, autocorrelation-based methods [15, 12, 17] are also not really direct measures of sample dynamics. It can be understood from the fact that the autocorrelation function and the power spectrum are tightly related by Wiener-Khinchin theorem. Note that, among the autocorrelation-based methods, OCDS [17] uses logarithmic intensity, and the Taylor series expansion of the logarithmic function also consists infinitely high order terms. In the other words, while logarithmic scaling helps reduce sensitivity to overall optical power fluctuations, it also amplifies local fluctuations and noise. And hence, OCDS is also not a direct measure of the sample dynamics.
Our new DOCT metrics, i.e., aLIV and Swiftness, are also not free from this issue. However, the numerical analyses revealed relatively straightforward relationships between these metrics and the sample dynamics (Section 3). Namely, aLIV mainly corresponds to DSR, while Swiftness mainly corresponds to the scatterer speed. These simple relationships enable easy interpretation of DOCT images. In addition, since the aLIV-DSR curves are nearly identical among all scatterer speeds, aLIV provides a more reliable estimate of the DSR than conventional LIV. This advantage is especially effective when analyzing real samples with a wide distribution of scatterer speeds. In addition, these properties of aLIV and Swiftness will also enable quantitative measurement of DSR and scatterer speed by combining with a sophisticated estimation theory as shown in a preliminary demonstration[47].